Akron Incorporated purchased an asset on September 1 of Year 1 for $375,000. The estimated residual value is $15,000. Akron estimates that the asset has a service life of 5 years. Calculate the depreciation for Year 1 using the sum-of-the-years'-digits method. $ Calculate the depreciation for Year 2 using the sum-of-the-years'-digits method.
Answer #1
Depreciation for Year 1 40000
Depreciation for Year 2 112000
Workings:
Sum-of-the-years'-digits 15 =5*(5+1)/2
Depreciation for Year 1 40000 =(375000-15000)*5/15*4/12
Depreciation for Year 2:
For remaining 1st year 80000 =(375000-15000)*5/15*8/12
For 2nd year 32000 =(375000-15000)*4/15*4/12
Depreciation for Year 2 112000